OpenStudy (anonymous):

Will medal Solve. -1/4y = 16

OpenStudy (matlee):

-64

OpenStudy (matlee):

idk

sammixboo (sammixboo):

First multiply -1 on both sides of the equation since you are dividing -1 in the orignal problem \(\Large \frac{-1}{4y}~=~16\) \(\Large -1 \times \frac{-1}{4y}~=~16\times -1\) \(\Large 4y~=~16\times -1\) So what is 16 x -1?

OpenStudy (anonymous):

-16

OpenStudy (studygurl14):

Um...@sammixboo you made a mistake

OpenStudy (studygurl14):

\(\large -1\times\frac{-1}{4y}\neq 4y\)

OpenStudy (here_to_help15):

dang again sammi tisk tisk tisk

OpenStudy (studygurl14):

Just do this: \(\large -\frac{1}{4}y=16\rightarrow -4(-\frac{1}{4}y)=16(-4)\rightarrow y = -64\)

OpenStudy (anonymous):

Thanks :)

sammixboo (sammixboo):

Good now we plug in -16 where 16 x -1 was \(\Large 4y~=~-16\) Now we must multiply 4 on both sides og the equation since we are dividing \(\Large 4y \times 4~=~-16\times4\) \(\Large y=~-64\)

OpenStudy (matlee):

Oh i see what you did there i didnt know you could do it like that

OpenStudy (studygurl14):

Strangely enough, although you made several mistakes @sammixboo , you still "got" the right answer.

OpenStudy (studygurl14):

You can't @matlee . Not really.

sammixboo (sammixboo):

I just skipped steps It cancels each other out and all

OpenStudy (matlee):

No she multiplied by 1 in the first place

sammixboo (sammixboo):

It's how I learned it

OpenStudy (matlee):

then she did the 4 after

OpenStudy (matlee):

although it was suppose to be 1/4 first then the 4

sammixboo (sammixboo):

I am not sure how it works, but it always turn out as the right answer. I never really got it

OpenStudy (studygurl14):

It defies the rules of algebra....

OpenStudy (studygurl14):

I think you made some mistakes, but you fixed it in your head, without showing every single step on here, so it came out okay. :) Weird, but if it works for you, then whatever. xD @sammixboo

sammixboo (sammixboo):

But I know your way is a right way I just never learned it like that

OpenStudy (here_to_help15):

Well Study no offense but at the end she was right and it seem to me like common sense so....

OpenStudy (matlee):

Im out to the next question

sammixboo (sammixboo):

:) I never been good at explaining math DX But my teacher just showed me another way to explain somewhat how I got my answers and that is how I have been doing it since I was 11

OpenStudy (studygurl14):

@Here_to_Help15 It's not common sense. Because... \(\large 4x\times 4=-16\times4\rightarrow16x=-64\)

sammixboo (sammixboo):

I think it is just some kind of cheat to skip some steps, but still get the same answer

OpenStudy (studygurl14):

(replace the x's with y's sorry)

sammixboo (sammixboo):

I know. It is that the 4x cancels out the times 4

OpenStudy (studygurl14):

Like I said, if you get the right answer, then by all means do it. :) Just cause I don't understand the process doesn't mean that it's a wrong process. Maybe one day you can teach me your method @sammixboo \(\large\ddot\smile\)

sammixboo (sammixboo):

My teacher explained a loooong explanation about how it works when I was younger, but I never really payed attention to that explanation, but it has always gotten me the right answer :)

OpenStudy (matlee):

Dont worry guys ill call my math teacher

OpenStudy (studygurl14):

:)